Can you help with this problem using a formula?
Find the sum of the even integers between -50 and 200 including them, that is, -50+(-48)+(-46)+...+198+200.
Thanks.

Answer by Alan3354(69443)

Can you help with this problem using a formula?
Find the sum of the even integers between -50 and 200 including them, that is, -50+(-48)+(-46)+...+198+200.
-------------------------
That's a total of 126 numbers. 1 + (200 - (-50))/2 = 126
The average is (200 + (-50))/2 = 75.
So the total is 126*75 = 9,450.